OpenStudy (anonymous):

im sorry I need a little help and it looks a lot like hollagirl24's. Because strangely where working on the same test. So here it is y=-x-7 y=2/3x+8

OpenStudy (compassionate):

System of equation?

OpenStudy (anonymous):

yeah

OpenStudy (compassionate):

Do you know the first step?

OpenStudy (anonymous):

no can you in numbers show me how to get the 1st step. Please I really need the help.

OpenStudy (compassionate):

y=-x-7 y=2/3x+8 So we're given two equation. You can both see they're solved for y, correct? So, doesn't that mean you can plug in numbers for y since you know what y is? In other words, if I had: x = 2 x = 2, you see, I already know that x = 1, so I can plug 1 into the other equation and say 2 = 2, which is a true statement

OpenStudy (compassionate):

Does that make sense?

OpenStudy (anonymous):

a little yeah were you trying to solve the x

OpenStudy (compassionate):

Yes. Let me give you another example y = x + 2 y = x + 3 So, I know what y equals so I can plug it in x + 2 = x + 3 You get what I'm doing?

OpenStudy (anonymous):

Okay I get so x=2

OpenStudy (compassionate):

Huh, did you just solve your original equation?

OpenStudy (anonymous):

what do you mean?

OpenStudy (compassionate):

\[-x-7 = \frac{ 2 }{ 3 }x + 8\]

OpenStudy (compassionate):

That's the first step. Now I want you to think about what I just did and why it makes sense. For example: y = 2x + 2 y = 2x + 1 You see. I KNOW y is equal to 2x + 2, so I can SUBSTITUTE the top equation into the bottom one 2x + 2 = 2x + 1

OpenStudy (compassionate):

Ready to move on to step 2?

OpenStudy (anonymous):

yes

OpenStudy (compassionate):

The next part is simplifying the expression. \[-x-7 = \frac{ 2 }{ 3 }x + 8 \rightarrow -x + x = \frac{ 2 }{ 3 } + 7 + 8 \rightarrow x = \frac{ 2 }{ 3 } + 15\] Can you add the last part, that is: \[ x = \frac{ 2 }{ 3 } + 15\]

OpenStudy (anonymous):

okay can you show me this using e problem that I sent in, You don't have to cheat. Just follow the steps to the problem leading to the answer without giving me the answer

OpenStudy (compassionate):

This IS the problem you posted...

OpenStudy (anonymous):

h

OpenStudy (anonymous):

I mean ok you can keep going

OpenStudy (compassionate):

(-9, 2) There, I just solved it for you.